# A theoretical analysis on the resolution of parametric PDEs via Neural Networks designed with Strassen algorithm

### v2

@@ 4 A Neural Network for Matrix Invertion
We know that for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}<1$ (recall Definition 2.5), we have the identity:
@@
The Neumann series can be used to approximate the inverse of matrices for which there exists $\alpha\in\mathbb{R}$ such that $\left\lVert\operatorname{I}-\alpha A\right\rVert_{2}<1$. In that case,
$A^{-1}=\alpha(\operatorname{I}_{n}-(\operatorname{I}_{n}-\alpha A))^{-1}=\alpha\sum_{k=0}^{\infty}(\operatorname{I}_{n}-\alpha A)^{k}.$
@@
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{DUP}_{n}}\right)\circ\dots\circ\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{DUP}_{n}}\right)}^{N}.$ (4.3)
@@
$\left\lVert A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (4.4)
@@
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2^{N-1}}}}\circ\Phi^{\operatorname{DUP}_{n}}\right)\circ\dots\circ\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2}}}\circ\Phi^{\operatorname{DUP}_{n}}\right)}^{N-1}\circ\Phi_{\frac{\varepsilon}{n},\frac{1}{2}}^{\operatorname{STR}_{n,\varrho}}\circ\Phi^{\operatorname{DUP}_{n}},$
it satisfies the same property. This can be seen by analyzing in detail the computations of the proof of Lemma 4.2. However, for clarity and coherence with the result needed for Theorem 4.5 below, we keep the subindices homogeneous and the second one equal to $1$.
@@
We can prove (4.4) by induction on $N$. We fix $A\in\mathbb{R}^{n\times n}$ with $\left\lVert A\right\rVert_{2}\leq\frac{1}{2}$.
@@
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{1}}\right)(A)-A^{2}\right\rVert_{2}$ $=\left\lVert R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)(A|A)-A^{2}\right\rVert_{2}$ (4.5)
@@
Suppose now that the statement (4.4) holds for $N\in\mathbb{N}$ and let us prove the case $N+1$. By (4.3), we know that:
$\Phi_{\varepsilon}^{\operatorname{SQR}^{N+1}_{n,\varrho}}=\Phi_{\varepsilon}^{\operatorname{SQR}^{1}_{n,\varrho}}\circ\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}.$
@@
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\leq\varepsilon+\left\lVert A\right\rVert_{2}^{2^{N}}\leq\varepsilon+\frac{1}{2^{2^{N}}}\leq\frac{1}{2}.$ (4.6)
@@
where we used the inductive hypothesis (4.4) and also used (4.5) since (4.6) holds. ∎
Remark 4.4.
This Matrix Neural Network can be used to compute any power of a matrix. Using the binary representation of the exponent, any power of a matrix can be expressed as the product of powers of 2 of that matrix. In other words, if we want to calculate $A^{M}$ with $A\in\mathbb{R}^{n\times n}$ and $M\in\mathbb{R}$, then, using the fact that there exist $N_{1},\dots,N_{k}\in\mathbb{N}\cup\left\{0\right\}$ all different such that $M=2^{N_{1}}+\dots+2^{N_{k}}$, $A^{M}=A^{2^{N_{1}}}\cdots A^{2^{N_{k}}}$. Since the product of matrices and the raising to a powers of 2 can be done with Matrix Neural Networks, the operation of raising a matrix to an arbitrary exponent can be expressed as such too.
Theorem 4.5.
@@
$\left\lVert\sum_{k=0}^{2^{N}-1}A^{k}-R\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (4.7)
for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$, satisyfing:
$\begin{split}M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)&\leq 14n^{\log_{2}7}(N-1)\left(M\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)+12\right)\\ &+n^{2}\left(L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)-14N+19+2\log_{2}n\right)+nN,\end{split}$ (4.8)
and satisfying:
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right].$ (4.9)
Also, if $N=1$ there exists a $\varrho$-MNN $\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}$ satisfying (4.7) for $\varepsilon=0$ and all $A\in\mathbb{R}^{n\times n}$, and satisfying:
$M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=n^{2}+n,$ (4.10)
$L\left(\Phi_{\varepsilon,K}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=1.$ (4.11)
Part of the improvements with respect to the paper [KPRS22] can be found in the Master Thesis [Rom24].
@@
  • Step 4: Proof of (4.7) for $N\geq 2$.
  • Step 5: Proof of (4.8) for $N\geq 2$.
  • Step 6: Proof of (4.9) for $N\geq 2$.
Step 1: The simple case $N=1$.
We define:
$\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}=((\operatorname{id},\operatorname{I}_{n},\operatorname{id})).$
Then, it is immediate that $R\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)(A)=A+\operatorname{I}_{n}$ with $M\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)=n^{2}+n$ and $L\left(\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}\right)=1$.
Step 2: Informal definition of the MNN.
We now suppose that $N\geq 2$.
The general idea is to use the product formula introduced in (4.2) for the Neumann series:
$\sum_{k=0}^{2^{N+1}-1}A^{k}=\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right).$
However, by our construction of $\Phi^{\operatorname{STR}}$, we require to bound the norm of the approximation of $\operatorname{I}_{n}+A^{2^{k}}$ and $A^{2^{k}}$ for $k\in\left\{0,\dots,N\right\}$ for an accurate product. For that purpose, we rewrite the previous product as follows:
$\begin{split}\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right)&=2^{\sum_{k=0}^{N}2^{k}}\prod_{k=0}^{N}2^{-2^{k}}\left(A^{2^{k}}+\operatorname{I}_{n}\right)\\ &=2^{2^{N+1}-1}\prod_{k=0}^{N}\left[\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right].\end{split}$ (4.12)
@@
The target MNN will then have two branches: one branch responsible for approximating $\left(\frac{A}{2}\right)^{2^{k}}$ for each $k\in\left\{0,\dots,N\right\}$ like in Lemma 4.2, while the other approximates the target product using recursively the different exponents of $\frac{A}{2}$. Our plan is then to follow the sketch:
$A$ $\mapsto\begin{pmatrix}\frac{A}{2}&|&\frac{A}{2}\\ \frac{A}{2}&|&0\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2}\\ \frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\end{pmatrix}$ (4.13)
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2}&|&\left(\frac{A}{2}\right)^{2}\\ \left(\frac{A}{2}\right)^{2}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2}&|&\frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{4}\\ \left(\frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\right)\left(\left(\frac{A}{2}\right)^{2}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2}\right)\end{pmatrix}\mapsto\cdots$
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2^{N-1}}&|&\left(\frac{A}{2}\right)^{2^{N-1}}\\ \left(\frac{A}{2}\right)^{2^{N-1}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{N-1}}&|&\prod_{k=0}^{N-2}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)\end{pmatrix}$
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2^{N}}\\ \prod_{k=0}^{N-1}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)\end{pmatrix}$
$\mapsto\left(\frac{A}{2}\right)^{2^{N}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{N}}\bigg{|}\prod_{k=0}^{N-1}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right)$
$\mapsto 2^{2^{N+1}-1}\prod_{k=0}^{N}\left(\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right).$
This way we compute $\left(\frac{A}{2}\right)^{2^{k}}$ only once for every $k\in\mathbb{N}$.
Step 3: Formal definition of the MNN.
We define some auxiliary MNN satisfying:
@@
$M\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=3n^{2},$ $L\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=1,$ (4.14)
@@
$\begin{cases}\Phi^{\operatorname{AUX}^{1}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{FILL}}_{n,L\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)}\right)\circ\Phi^{\operatorname{DUP}}_{n},\\ \Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\circ\Phi^{\operatorname{MIX}}_{n,i-1}\circ\Phi^{\operatorname{AUX}^{i-1}_{n,\varrho}}_{\varepsilon}&\forall i\geq 2.\end{cases}$ (4.15)
@@
$\left((\mathcal{L}^{1},C^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L})\right)\coloneq\Phi^{\operatorname{STR}_{n,\varrho}}_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}\circ\Phi^{\operatorname{FLIP}}_{n,N-1}\circ\Phi^{\operatorname{AUX}^{N-1}_{n,\varrho}}_{2^{1-2^{N}}\varepsilon},$
@@
$\operatorname{SQR}_{\varepsilon}^{i}(A)$ $=\left[R\left(\Phi_{\varepsilon}^{\operatorname{AUX}^{i}_{n,\varrho}}\right)(A)\right]_{1:n,1:n},$
@@
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$. It is not difficult to see using (4.3) and (4.15) that:
$\operatorname{SQR}^{i}_{\varepsilon}(A)=R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}\right)\left(\frac{A}{2}\right),$ (4.16)
where $\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}$ is the $\varrho$-MNN from Lemma 4.2. Thus, we know by (4.4) that:
$\left\lVert\operatorname{SQR}^{i}_{\varepsilon}(A)\right\rVert_{2}\leq\varepsilon+\left\lVert\frac{A}{2}\right\rVert_{2}^{2^{i}}\leq\varepsilon+2^{-2^{i}}$ (4.17)
@@
$\left\lVert P^{i}(A)\right\rVert_{2}\leq\prod_{k=0}^{i}\frac{2}{2^{2^{k}}}=2^{i+2-2^{i+1}}\leq\frac{1}{2}$ (4.18)

### v3

@@ 3.2 Matrix Inversion Approximation
We know that for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}<1$, we have the identity:
@@
The Neumann series can be used to approximate the inverse of any matrix since a matrix $A$ is invertible if and only if there is a (invertible) matrix $P$ for which $\left\lVert\operatorname{I}-AP\right\rVert_{2}<1$. In that case,
$A^{-1}=P(\operatorname{I}_{n}-(\operatorname{I}_{n}-AP))^{-1}=P\sum_{k=0}^{\infty}(\operatorname{I}_{n}-AP)^{k}.$
The particular case when $P=\alpha\operatorname{I}$ is of high interest for symmetric positive definite matrices, as $\alpha$ can be estimated fairly easily.
We can finally get a MNN approximating the inverse of some matrices.
Theorem 3.7.
$N(\varepsilon,\delta)$ $=\max\left\{\left\lceil\log_{2}\left(\frac{\log_{2}(\varepsilon(1-\delta))}{\log_{2}(\delta)}\right)\right\rceil,1\right\},$ (3.20)
$\Sigma(\varepsilon,\delta,n)$ $=2^{-2^{N(\varepsilon,\delta)}}\frac{\min\left\{\varepsilon,1/4\right\}}{16n^{3}}.$
for all $A\in\mathbb{R}^{n\times n}$ such that
and satisfying:
$M\left(\Phi^{\operatorname{INV}_{n,\varrho}^{\alpha}}_{\varepsilon,\delta}\right)$ $\leq 14n^{\log_{2}7}\left(N\left(\frac{\varepsilon}{2\alpha},\delta\right)-1\right)\left(M\left(\Pi^{\varrho}_{\Sigma(\varepsilon/\alpha,\delta,n),2n}\right)+12\right)$
$+n^{2}\left(L\left(\Pi_{\Sigma(\varepsilon/\alpha,\delta,n),2n}^{\varrho}\right)-14N\left(\frac{\varepsilon}{2\alpha},\delta\right)+20+2\log_{2}n\right)$ (3.22)
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}^{\alpha}_{n,\varrho}}\right)$ $\leq N\left(\frac{\varepsilon}{2\alpha},\delta\right)\left[2\log_{2}n+5+L\left(\Pi_{\Sigma(\varepsilon/\alpha,\delta,n),2n}^{\varrho}\right)\right],$ (3.23)
if $N\left(\frac{\varepsilon}{2\alpha},\delta\right)=1$.
Note that $\varepsilon\mapsto N(\varepsilon,\delta)$ is a decreasing function when $\delta\in(0,1)$.
The proof can be found in Appendix B. The idea of the proof is first construct a MNN that approximates the powers $A^{2^{k}}$, then another responsible of the product (3.19) and finally, using the Neumann series for an appropriate number of terms ($2^{N(\frac{\varepsilon}{2\alpha},\delta)+1}$ concretely), we get the desired MNN.
Remark 3.8.
As stated in Remark 3.6, we could construct a $\varrho$-MNN that approximates any matrix $A\in\mathbb{R}^{d\times d}$ such that
$\left\lVert\operatorname{I}_{d}-AP\right\rVert\leq\delta<1,$
where is $P\in\mathbb{R}^{d\times d}$ is a fixed invertible matrix. But for simplicity of the statement and proof, we stick to the case $P=\alpha\operatorname{I}_{d}$.
Remark 3.9.
For the particular case $\varrho=\operatorname{ReLU}$ and $\alpha=1$, estimates (3.22) and (3.23) can be written as
$L\left(\Phi_{\varepsilon,\delta}^{\operatorname{INV}_{n}^{*}}\right)$ $\leq C\log_{2}m(\varepsilon,\delta)\left(\log_{2}\frac{1}{\varepsilon}+m(\varepsilon,\delta)+\log_{2}n\right),$ (3.27)
where $C>0$ is a constant and
$m(\varepsilon,\delta)=\frac{\log_{2}(\varepsilon(1-\delta)/2)}{\log_{2}\delta}\approx 2^{N(\varepsilon/2,\delta)}.$
This result presents various improvements with respect to prior work, concretely [kutyniok2022theoretical, Theorem 3.8]. Although the number of layers is asymptotically worse, the upper bound on the number of weights on (3.26) is an improvement over theirs. The exponent on the dimension of the input matrices is reduced from 3 to $\log_{2}7$ and the factor is $\log_{2}m(\varepsilon,\delta)$ instead of $m(\varepsilon,\delta)\log_{2}^{2}(m(\varepsilon,\delta))$. In parentheses, even if we have $m(\varepsilon,\delta)$ instead of $\log_{2}m(\varepsilon,\delta)$, it is still an improvement if we compare the expanded expression $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))$ to $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))\log_{2}^{2}(m(\varepsilon,\delta))$. The first change comes from the use of the Strassen algorithm for matrix multiplication, while the second is due to the use of the identity (3.19).
Remark 3.10.
In [kutyniok2022theoretical], they consider parametric coercive PDEs whose discretizations, by applying the Galerkin method, results in a linear system whose respective matrices are symmetric positive definite. They prove that the inverse of all of those matrices can be approximated by a single Neural Network (with worse upper bounds on the number of parameters as explained in Remark 3.9). Then, we reach the same conclusion with a better dependency on the Galerkin dimension for approximating the parameter-to-solution map. However, other methods, like those of multilevel ideas in [JMLR:v24:23-0421], provides even better bounds for the particular equation considered where they relay in a more efficient way of solving the resulting linear system.
4 Conclusions
We have constructed two important Neural Networks approximating matrix multiplication and inversion operators. We have been able to do so without restricting ourselves to a fixed activation function. Also, we bound their number of weights and layers, and by using the Strassen algorithm we have reduced a potential cubic dependency on the dimension to a $\log_{2}7$ exponent. Rewriting the Neumann partial sum, we reduce the upper bounds on the number of parameters by reducing the number of matrix multiplication apparitions.
Appendix A Proof of Proposition 3.1
Proof of 3.1.
Let us suppose that Proposition 3.1 is true for k1k-1. Let us start by showing (3.4). Using the Strassen algorithm, we know that (2.2) is satisfied. Consequently, because of the triangular inequality, defining $P_{i}$ as in (2.3), it suffices to show that:
[…]
Let us show the estimate for $i=1$, as the other six estimates are analogous. Since $\left\lvert A\right\rvert_{\infty}\leq K$, then $\left\lvert A[1,1]+A[2,2]\right\rvert_{\infty}\leq 2K$. Similarly, since $\left\lvert B\right\rvert_{\infty}\leq K$, we have $\left\lvert B[1,1]+B[2,2]\right\rvert_{\infty}\leq 2K$. Thus, using the recursive hypothesis with k1k-1 we obtain that:
[…]
concluding the proof of (3.4).
Using the previous equality, (3.2) and (3.3), we obtain recursively:
$M\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)$ $=M\left(\Phi_{\varepsilon,K}^{\operatorname{PAR}_{k,\varrho}}\right)+M\left(\Phi^{\operatorname{MIX}_{k}}\right)+M\left(\Phi^{\operatorname{SPLIT}_{k}}\right)$ (A.2)
$=\cdots$
We now calculate the sum using the fact that it is geometric:
$9\sum_{i=1}^{k}4^{i}\cdot 7^{k-i}=9\cdot 7^{k}\sum_{i=1}^{k}\left(\frac{4}{7}\right)^{i}=9\cdot 7^{k}\frac{4/7-(4/7)^{k+1}}{1-4/7}=12\left(7^{k}-4^{k}\right).$ (A.3)
Putting together (A.2) and (A.3), we obtain (3.5).
Thus, using induction, we conclude with (3.6). ∎
Appendix B Proof of Theorem 3.7
@@ 4 Conclusions
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\odot\Phi^{\operatorname{DUP}_{n}}\right)\odot\dots\odot\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\odot\Phi^{\operatorname{DUP}_{n}}\right)}^{N}.$ (B.1)
@@
$\left\lVert A^{2^{N}}-R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (B.2)
@@
$\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{N}}=\overbrace{\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2^{N-1}}}}\odot\Phi^{\operatorname{DUP}_{n}}\right)\odot\dots\odot\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},\varepsilon+\frac{1}{2^{2}}}\odot\Phi^{\operatorname{DUP}_{n}}\right)}^{N-1}\odot\Phi_{\frac{\varepsilon}{n},\frac{1}{2}}^{\operatorname{STR}_{n,\varrho}}\odot\Phi^{\operatorname{DUP}_{n}},$
it satisfies the same property. This can be seen by analyzing in detail the computations of the proof of Lemma B.1. However, for clarity and coherence with the result needed for Lemma B.3 below, we keep the subindices homogeneous and the second one equal to $1$.
@@
We can prove (B.2) by induction on $N$. We fix $A\in\mathbb{R}^{n\times n}$ with $\left\lVert A\right\rVert_{2}\leq\frac{1}{2}$.
@@
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}_{n,\varrho}^{1}}\right)(A)-A^{2}\right\rVert_{2}$ $=\left\lVert R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)(A|A)-A^{2}\right\rVert_{2}$ (B.3)
@@
Suppose now that the statement (B.2) holds for $N\in\mathbb{N}$ and let us prove the case $N+1$. By (B.1), we know that:
$\Phi_{\varepsilon}^{\operatorname{SQR}^{N+1}_{n,\varrho}}=\Phi_{\varepsilon}^{\operatorname{SQR}^{1}_{n,\varrho}}\odot\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}.$
@@
$\left\lVert R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{N}_{n,\varrho}}\right)(A)\right\rVert_{2}\leq\varepsilon+\left\lVert A\right\rVert_{2}^{2^{N}}\leq\varepsilon+\frac{1}{2^{2^{N}}}\leq\frac{1}{2}.$ (B.4)
@@
where we used the inductive hypothesis (B.2) and also used (B.3) since (B.4) holds. ∎
Now, we construct a MNN that approximates the product (3.19).
Lemma B.3.
@@
$\left\lVert\sum_{k=0}^{2^{N}-1}A^{k}-R\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)(A)\right\rVert_{2}\leq\varepsilon$ (B.5)
for all $A\in\mathbb{R}^{n\times n}$ such that $\left\lVert A\right\rVert_{2}\leq 1$, satisfying:
$\begin{split}M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)&\leq 14n^{\log_{2}7}(N-1)\left(M\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)+12\right)\\ &+n^{2}\left(L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)-14N+19+2\log_{2}n\right)+nN,\end{split}$ (B.6)
and
$L\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}\right)\leq N\left[2\log_{2}n+5+L\left(\Pi_{2^{-2^{N}}\frac{\varepsilon}{8n^{3}},2n}^{\varrho}\right)\right].$ (B.7)
Also, if $N=1$ there exists a $\varrho$-MNN $\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{N}}$ satisfying (B.5) for $\varepsilon=0$ and all $A\in\mathbb{R}^{n\times n}$, and satisfying:
$M\left(\Phi_{\varepsilon}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=n^{2}+n,$ (B.8)
$L\left(\Phi_{\varepsilon,K}^{\operatorname{NEU}_{n,\varrho}^{1}}\right)$ $=1.$ (B.9)
Part of the improvements with respect to the paper [kutyniok2022theoretical] can be found in the Master Thesis [romera2024neuralnetworksnumericalanalysis]
@@
  • Step 4: Proof of (B.5) for $N\geq 2$.
  • Step 5: Proof of (B.6) for $N\geq 2$.
  • Step 6: Proof of (B.7) for $N\geq 2$.
Step 1: The simple case $N=1$. We just have define $\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}=((\operatorname{id},\operatorname{I}_{n},\operatorname{id}))$, from where the result follows easily.
Step 2: Informal definition of the MNN. We now suppose that $N\geq 2$. The general idea is to use the product formula introduced in (3.19) for the Neumann series. However, by our construction of $\Phi^{\operatorname{STR}}$, we require to bound the norm of the approximation of $\operatorname{I}_{n}+A^{2^{k}}$ and $A^{2^{k}}$ for $k\in\left\{0,\dots,N\right\}$ for an accurate product. For that purpose, we rewrite (3.19) as follows:
$\prod_{k=0}^{N}\left(A^{2^{k}}+\operatorname{I}_{n}\right)=2^{\sum_{k=0}^{N}2^{k}}\prod_{k=0}^{N}2^{-2^{k}}\left(A^{2^{k}}+\operatorname{I}_{n}\right)=2^{2^{N+1}-1}\prod_{k=0}^{N}\left[\left(\frac{A}{2}\right)^{2^{k}}+\left(\frac{\operatorname{I}_{n}}{2}\right)^{2^{k}}\right].$ (B.10)
@@
The target MNN will then have two branches: one branch responsible for approximating $\left(\frac{A}{2}\right)^{2^{k}}$ for each $k\in\left\{0,\dots,N\right\}$ like in Lemma B.1, while the other approximates the target product using recursively the different exponents of $\frac{A}{2}$. Our plan is then to follow the sketch:
$A$ $\mapsto\begin{pmatrix}\frac{A}{2}&|&\frac{A}{2}\\ \frac{A}{2}&|&0\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2}\\ \frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\end{pmatrix}\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{4}\\ \left(\frac{A}{2}+\frac{\operatorname{I}_{n}}{2}\right)\left(\left(\frac{A}{2}\right)^{2}+\frac{1}{2^{2}}\operatorname{I}_{n}\right)\end{pmatrix}\mapsto\cdots$ (B.11)
$\mapsto\begin{pmatrix}\left(\frac{A}{2}\right)^{2^{N}}\\ \prod_{k=0}^{N-1}\left(\left(\frac{A}{2}\right)^{2^{k}}+\frac{1}{2^{2^{k}}}\operatorname{I}_{n}\right)\end{pmatrix}\mapsto 2^{2^{N+1}-1}\prod_{k=0}^{N}\left(\left(\frac{A}{2}\right)^{2^{k}}+\frac{1}{2^{2^{k}}}\operatorname{I}_{n}\right).$
In this way, we compute $\left(\frac{A}{2}\right)^{2^{k}}$ only once for every $k\in\mathbb{N}$.
Step 3: Formal definition of the MNN. We define some auxiliary MNN satisfying:
@@
$M\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=3n^{2},$ $L\left(\Phi^{\operatorname{DUP}}_{n}\right)$ $=1,$ (B.12)
@@
$\begin{cases}\Phi^{\operatorname{AUX}^{1}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{FILL}}_{n,L\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)}\right)\odot\Phi^{\operatorname{DUP}}_{n},\\ \Phi^{\operatorname{AUX}^{i}_{n,\varrho}}_{\varepsilon}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1},\Phi^{\operatorname{STR}_{n,\varrho}}_{\frac{\varepsilon}{4n},1}\right)\odot\Phi^{\operatorname{MIX}}_{n,i-1}\odot\Phi^{\operatorname{AUX}^{i-1}_{n,\varrho}}_{\varepsilon}&\forall i\geq 2.\end{cases}$ (B.13)
@@
$\left((\mathcal{L}^{1},C^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L})\right)\coloneq\Phi^{\operatorname{STR}_{n,\varrho}}_{2^{1-2^{N}}\frac{\varepsilon}{4n},1}\odot\Phi^{\operatorname{FLIP}}_{n,N-1}\odot\Phi^{\operatorname{AUX}^{N-1}_{n,\varrho}}_{2^{1-2^{N}}\varepsilon},$
@@
$\operatorname{SQR}_{\varepsilon}^{i}(A)$ $=\left[R\left(\Phi_{\varepsilon}^{\operatorname{AUX}^{i}_{n,\varrho}}\right)(A)\right]_{1:n,1:n},$ (B.14)
@@
for all $i\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$. It is not difficult to see using (B.1) and (B.13) that:
$\operatorname{SQR}^{i}_{\varepsilon}(A)=R\left(\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}\right)\left(\frac{A}{2}\right),$ (B.15)
where $\Phi_{\varepsilon}^{\operatorname{SQR}^{i}_{n,\varrho}}$ is the $\varrho$-MNN from Lemma B.1. Thus, we know by (B.2) that:
$\left\lVert\operatorname{SQR}^{i}_{\varepsilon}(A)\right\rVert_{2}\leq\varepsilon+\left\lVert\frac{A}{2}\right\rVert_{2}^{2^{i}}\leq\varepsilon+2^{-2^{i}}$ (B.16)
@@
$\left\lVert P^{i}(A)\right\rVert_{2}\leq\prod_{k=0}^{i}\frac{2}{2^{2^{k}}}=2^{i+2-2^{i+1}}\leq\frac{1}{2}$ (B.17)